OpenStudy (happinessbreaksbones):

help.. 2x + 4 = –3x + 14

OpenStudy (happinessbreaksbones):

I think I know the answer but I want to make sure

OpenStudy (anonymous):

tell your answer

OpenStudy (happinessbreaksbones):

x = 2

OpenStudy (anonymous):

you're right (:

OpenStudy (happinessbreaksbones):

HAH well thanks :)

OpenStudy (anonymous):

Ok ! See: 5x+4=14 5x=14-4 5x=10 x=10/5 X=2

OpenStudy (happinessbreaksbones):

thank you :) thanks everyone!